OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x3 + x2 - 11

OpenStudy (anonymous):

Having trouble with this. The relative is deferent then the local correct? Sorry to bother you again, if your still up, have any idea on this one? @satellite73 @Luigi0210

OpenStudy (anonymous):

has a min, not max since it is a parabola that opens up min is the second coordinate of the vertex

OpenStudy (anonymous):

oh crap it is cubic!! sorry

OpenStudy (anonymous):

\[f(x) = 2x^3 + x^2 - 11\] \[f'(x)=3x^2+2x\] set it equal to zero and solve it has a relative max and a relative min

OpenStudy (anonymous):

it is cubic so it goes from \(-\infty\) to \(\infty\) and has no absolute max or min

OpenStudy (anonymous):

if you solve \[3x^2+2x=0\] you get \(x=-\frac{2}{3}\) or \(x=0\)

OpenStudy (anonymous):

the one on the left will give a relative max and \(0\) will give a relative min

OpenStudy (anonymous):

answers

OpenStudy (anonymous):

oh and no relative is local same thing

OpenStudy (anonymous):

oh damn i made a mistake

OpenStudy (anonymous):

\[f'(x)=6x^2+2x\]

OpenStudy (anonymous):

but none of your answer are answers to this one

OpenStudy (anonymous):

I mean its from my high school I think there has to be one

OpenStudy (anonymous):

check the function again, i am quite positive these are not answers to that one

OpenStudy (anonymous):

f(x) = 2x^3 + x^2 - 11

OpenStudy (anonymous):

did you know that?

OpenStudy (anonymous):

\[f(x) = 2x^3 + x^2 - 11\] \[f'(x)=6x^2+2x\] for sure now set \(6x^2+2x=0\) you get \[2x(3x+1)=0\] so either \(x=0\) or \(x=-\frac{1}{3}\)

OpenStudy (anonymous):

your answers are not to this question, or someone made a mistake sorry

OpenStudy (anonymous):

hmm interesting. if you had to make a guess?

OpenStudy (anonymous):

they are not even close don't guess, they are wrong

OpenStudy (anonymous):

well its for online highschool and if i dont check any then its definitely wrong. Heres a pic of the whole question. Maybe you miss understood something...

OpenStudy (anonymous):

yeah i see it it is wrong

OpenStudy (anonymous):

relative min is \(-11\) at \(x=0\)

OpenStudy (anonymous):

look i know this is on line and you have to put in something, but you should contact them somehow and tell them it is wrong don't give in to the machine anyway, if you have a real teacher he or she will think you are paying attention and be happy

OpenStudy (anonymous):

the answers are not even close

OpenStudy (anonymous):

okay i sent a message in. Thanks

OpenStudy (anonymous):

I saw you answered a similar question. What polynomial has a graph that passes through the given points? (-2, 2) (-1, -1) (1, 5) (3, 67)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Mathard